3^/x - 5 = 3........... 3 cuberoot over x - 5 = 3
posted my question incorrectly
Thanks

Answer by stanbon(75887)

3^/x - 5 = 3
-----------
x^(1/3) - 5 = 3
x^(1/3) = 8
x = 8^3
x = 512
===========
Cheers,
Stan H.